OpenStudy (anonymous):

A wild birdseed mix is made by combining 100lb of millet seed costing $.60 per pound with sunflower seeds costing $1.10 per pound. How many pounds of sunflower seeds are needed to make a mixture that costs $.70 per pound?

OpenStudy (btaylor):

create a system: m + s = 110 --> m = 110 - s 0.6m + 1.1s = (100)(0.70) now substitute the value for m in the second equation: 0.6(110 - s) + 1.1s = 70 66-0.6s+1.1s=70 0.5s=4 s=8 --> m = 92

OpenStudy (anonymous):

Thanks taylor my answer is difinitely wrong. IT is 40 thanks for teaching me.